OpenStudy (anonymous):

In a math class of 26 students each girl drew a triangle and each boy drew a rectangle if there were 92 sides drawn by students how many girls were in the class

OpenStudy (anonymous):

12 girls

OpenStudy (anonymous):

girl : boy -student 3 : 4 - 1student 6 : 8 - 2students 9 : 12 - 3students ... ... 36 : 48 -12students 39 : 52 -13students 42 : 56 -14students *take 36 for girl and 56 for boy 36+56=92 sides so; 36 side were drawn by 12 girls

OpenStudy (anonymous):

You can also solve by using two equations and substitution. 3G + 4B = 92 G + B = 26